# Non-equilibrium lifetimes of DNA under electronic current in a molecular junction

### v1

@@ I Introduction
It is well established that electric fields and currents can influence the structure of the molecule in an electronic junction, from field-induced chemical reactions to current-driven conformational switching [2, 3, 4]. We propose the microscopic mechanisms by which tunnelling electrons destabilise – or in some cases stabilise – DNA in a molecular junction environment. Quantum electronic current couples to classical mechanical motion of the DNA. Such coupling injects significant energy into mechanical motion, modifies the potential energy surface, and generates spatially varying fluctuations and dissipation, leading to complex and sometimes counterintuitive effects.
@@ II.1 Hamiltonian
The electrodes are modeled as non-interacting electron reservoirs:
@@
Here, $j$ indexes base pairs along the chain, while $\beta\in{A,B}$ labels the DNA strands. The on-site energies are denoted by $\epsilon_{\beta}$, with intra-strand hopping integrals $t_{\beta}$ and inter-strand hopping $v$. Molecular parameters are the same as in our previous paper [10]: $\epsilon_{A}=-0.2$ eV, $\epsilon_{B}=0.1$ eV, $t_{A}=0.1$ eV, $t_{B}=-0.14$ eV, and v=0.08v=-0.08 eV.
@@ II.2 Green’s functions and self-energies
The retarded component of leads self-energy is computed using wide-band approximation
@@
$\Gamma_{\alpha;\beta i\sigma,\beta^{\prime}i^{\prime}\sigma^{\prime}}=2\pi\delta_{\beta\beta^{\prime}}\delta_{\sigma\sigma^{\prime}}\sum_{k}t^{*}_{\alpha k,\beta i\sigma}\delta(\omega-\epsilon_{k\alpha})t_{\alpha k,\beta i^{\prime}\sigma},$ (11)
and assumed to be $\omega$-independent. We use $\Gamma_{\alpha;\beta i\sigma,\beta^{\prime}i^{\prime}\sigma^{\prime}}=0.1$ eV in our calculations.
The lesser and greater self-energies are computed as
@@
All components of the Green’s functions depend instantaneously on the DNA’s mechanical degree of freedom and are defined using standard adiabatic relations:
@@ II.3 Keldysh-Langevin molecular dynamics
The mechanical degree of freedom of DNA is considered a classical variable within our approach. If we additionally assume that DNA’s conformational dynamics are slow relative to the tunnelling electrons, we can obtain a Langevin equation of motion [13, 14, 15, 16, 17, 18, 6, 5]:
@@ III.1 Electric current reshapes effective potential energy surface for DNA’s mechanical motion
where the electronic force, denoted as $F_{\text{e,neq}}$, is given by (20) and the choice of $x_{0}$ is arbitrary. The effective potential energy contains the equilibrium part, which is the classical potential from (6), and a modification to the potential due to current-induced force.
@@
Our subsequent analysis of DNA mechanical stability requires that we select the threshold value for mechanical motion which marks the onset of DNA’s structural instability - we assume once this critical value is reached, the DNA experiences structural instability resulting in denaturation. The critical threshold ($x_{b}=1$ a.u.) is selected. This choice is somewhat arbitrary given the collective nature of the DNA’s mechanical coordinate in our model but can be physically motivated by typical ranges of realistic base-pairing potentials in DNA. The threshold $x_{b}=1$ a.u. has a corresponding critical potential energy $U(x_{b})$ which will be treated as the energy barrier $U_{\text{bar}}$ that the mechanical degree of freedom has to overcome to reach an unstable configuration.
@@ III.2 DNA non-equilibrium lifetime as mean first passage time
We define first-passage time as the time it takes for the DNA’s mechanical coordinate, evolving according to the generalized Langevin equation, to reach a predefined critical displacement $x_{b}$ for the first time. Starting from an initial state sampled from the stationary distribution, the trajectory is propagated under the influence of deterministic forces (mechanical restoring force, non-equilibrium electronic force), viscosity (environmental and electronic), and stochastic forces (environmental and electronic noise). The first-passage time $\tau$ is then recorded as the elapsed simulation time until the coordinate $x(t)$ crosses the critical threshold $x_{b}$ (taken here as $x_{b}=1$ a.u.) for the first time:
@@
III.3 Landauer blowtorch effect emerges from localized electronic fluctuating and dissipating forces
The Landauer blowtorch effect refers to the kinetic destabilisation/stabilisation of a system when local fluctuations are spatially inhomogeneous. Even if the average potential landscape favours stability, regions of enhanced local noise can effectively lower barriers and promote escape, as if a “blowtorch” were applied selectively along the reaction coordinate. The reverse is also observed at low voltages where regions of high temperature and moderate viscosity lead to diminished local density, which if located at the critical point can impede escapes and, thus, enhance stability. Within the Langevin framework, this is captured by a coordinate-dependent effective temperature $T(x)$ defined through the generalized fluctuation-dissipation relation [21]
$T(x)=\frac{D_{e}(x)+D_{\text{env}}}{2(\xi_{e}(x)+\xi_{\text{env}})},$ (29)
The presence of non-equilibrium electronic viscosity and dissipation results into inhomogeneous effective temperature $T(x)$. When $T(x)$ varies significantly with $x$, the system no longer behaves as having a single temperature, leading to locally biased fluctuations and dissipation.
@@ III.3 Landauer blowtorch effect emerges from localized electronic fluctuating and dissipating forces
Fig. 3(c) compares trajectory probability densities with and without the inclusion of electronic viscosity and noise; the figure also shows the spatial dependence of the effective temperature for $\chi=10$ meV at $V=1$ V. When only the electron force and environmental fluctuation–dissipation are included (orange dashed), the distribution is narrowly localized around the potential minimum. By contrast, the full treatment including electronic viscosity and noise (blue dashed) broadens the distribution, enabling access to higher-energy displacements even in the presence of a confining potential. This is a direct signature of the blowtorch effect: non-uniform dissipation and stochastic forces effectively create spatial variations in “temperature”, making escape pathways more accessible.

### v2

@@ I Introduction
It is well established that electric fields and currents can influence the structure of the molecule in an electronic junction, from field-induced chemical reactions to current-driven conformational switching [2, 3, 4, 5]. We propose the microscopic mechanisms by which tunnelling electrons destabilise – or in some cases stabilise – DNA in a molecular junction environment. Quantum electronic current couples to classical mechanical motion of the DNA. Such coupling injects significant energy into mechanical motion, modifies the potential energy surface, and generates spatially varying fluctuations and dissipation, leading to complex and sometimes counterintuitive effects.
@@ II.1 Hamiltonian
The electrodes are modelled as non-interacting electron reservoirs:
@@
Here, $j$ indexes base pairs along the chain, while $\beta\in{A,B}$ labels the DNA strands. The on-site energies are denoted by $\epsilon_{\beta}$, with intra-strand hopping integrals $t_{\beta}$ and inter-strand hopping $v$. Molecular parameters are the same as in our previous paper [11]: $\epsilon_{A}=-0.2$ eV, $\epsilon_{B}=0.1$ eV, $t_{A}=0.1$ eV, $t_{B}=-0.14$ eV, and v=0.08v=-0.08 eV. The number of DNA basepairs: $N=10$.
@@ II.2 Green’s functions and self-energies
The retarded component of leads self-energy is computed using wide-band approximation [14]
@@
$\Gamma_{\alpha;\beta i\sigma,\beta^{\prime}i^{\prime}\sigma^{\prime}}=2\pi\delta_{\beta\beta^{\prime}}\delta_{ii^{\prime}}\delta_{\sigma\sigma^{\prime}}\sum_{k}t^{*}_{\alpha k,\beta i\sigma}\delta(\omega-\epsilon_{k\alpha})t_{\alpha k,\beta i\sigma},$ (11)
and assumed to be $\omega$-independent. Matrix $\Gamma_{\alpha;\beta i\sigma,\beta^{\prime}i^{\prime}\sigma^{\prime}}$ is diagonal in strand ($\beta$), spin ($\sigma$), and nucleotide ($i$) indices and is nonzero only for terminal nucleotides ($i=1,N$) directly coupled to the electrodes. This represents local, spin-independent coupling between the electrode states and the outermost base pairs of each strand. We use $\Gamma_{\alpha;\beta 1\sigma,\beta 1\sigma}=\Gamma_{\alpha;\beta N\sigma,\beta N\sigma}=0.1$ eV in our calculations.
The lesser and greater self-energies are computed as [14]
@@
All components of the Green’s functions depend instantaneously on the DNA’s mechanical degree of freedom and are defined using standard adiabatic relations [14]:
@@ II.3 Keldysh-Langevin molecular dynamics
The mechanical degree of freedom of DNA is considered a classical variable within our approach. Quantum vibrational effects play an important role in inelastic electron transport through molecules [14]. In this work, however, the collective coordinate represents a slow, large-scale conformational motion of the DNA molecule , whose characteristic frequency (5 meV) is much smaller than room temperature. In this limit, the motion is well described classically. If we additionally assume that DNA’s conformational dynamics are slow relative to the tunnelling electrons, we can obtain a Langevin equation of motion [15, 16, 17, 18, 19, 20, 7, 6]:
@@ III.1 Electric current reshapes effective potential energy surface for DNA’s mechanical motion
where the electronic force, denoted as $F_{\text{e,neq}}$, is given by (20) and the choice of $x_{0}$ is arbitrary. The effective potential energy contains the equilibrium part, which is the classical potential from (6), and a modification to the potential due to current-induced force. The effective potential $U(x)$ represents only the conservative component of the electronic force. It provides intuitive insight into how the electronic current reshapes the deterministic part of the energy landscape. However, it is critical to include non-conservative viscous and stochastic forces, which are also explicitly computed in our Keldysh-Langevin molecular dynamics.
@@
Our subsequent analysis of DNA mechanical stability requires that we select the threshold value for mechanical motion which marks the onset of DNA’s structural instability - we assume once this critical value is reached, the DNA experiences structural instability. The critical threshold ($x_{b}=1$ a.u.) is selected. Molecular dynamics studies of base-pair deformability in double stranded DNA consistently show that conformational displacements of order 1-5 a.u of nucleotides from Watson-Crick structure correspond to the onset of DNA mechanical instability [23, 24, 25]. In our coarse-grained model, the coordinate $x$ represents a collective conformational mode (bending, twisting, or stretching of multiple base pairs). A displacement of 1 a.u. along this collective coordinate therefore corresponds to the lower bound of the physical range for destabilisation of individual base pairs seen in atomistic studies.
This choice is somewhat arbitrary given the collective nature of the DNA’s mechanical coordinate in our model but can be physically motivated by typical ranges of realistic base-pairing potentials in DNA. The threshold $x_{b}=1$ a.u. has a corresponding critical potential energy $U(x_{b})$ which will be treated as the energy barrier $U_{\text{bar}}$ that the mechanical degree of freedom has to overcome to reach an unstable configuration.
@@ III.2 DNA non-equilibrium lifetime as mean first passage time
We define first-passage time as the time it takes for the DNA’s mechanical coordinate, evolving according to the generalised Langevin equation, to reach a predefined critical displacement $x_{b}$ for the first time. Starting from an initial state sampled from the stationary distribution, the trajectory is propagated under the influence of deterministic forces (mechanical restoring force, non-equilibrium electronic force), viscosity (environmental and electronic), and stochastic forces (environmental and electronic noise). The first-passage time $\tau$ is then recorded as the elapsed simulation time until the coordinate $x(t)$ crosses the critical threshold $x_{b}$ (taken here as $x_{b}=1$ a.u.) for the first time:
@@
III.3 Landauer blowtorch effect emerges from localised electronic fluctuating and dissipating forces
The Landauer blowtorch effect refers to the kinetic destabilisation/stabilisation of a system when local fluctuations are spatially inhomogeneous. Even if the average potential landscape favours stability, regions of enhanced local noise can effectively lower barriers and promote escape, as if a “blowtorch” were applied selectively along the reaction coordinate. The reverse is also observed at low voltages where regions of high temperature and moderate viscosity lead to diminished local density, which if located at the critical point can impede escapes and, thus, enhance stability. Within the Langevin framework, this is captured by a coordinate-dependent effective temperature $T(x)$ defined through the generalised fluctuation-dissipation relation [26]
$T(x)=\frac{D_{e}(x)+D_{\text{env}}}{2(\xi_{e}(x)+\xi_{\text{env}})}.$ (29)
Eq.(29) provides a definition of the local effective temperature $T(x)$ rather than a true fluctuation-dissipation theorem. Only in equilibrium - that is, when the left and right chemical potentials are equal - does Eq.(29) reduce to the genuine fluctuation-dissipation relation. In this limit, the $x$-dependences of the numerator and denominator cancel, and $T(x)$ coincides with the electrode temperature. The presence of non-equilibrium electronic viscosity and random forces result into inhomogeneous effective temperature $T(x)$. When $T(x)$ varies significantly with $x$, the system no longer behaves as having a single temperature, leading to locally biased fluctuations and dissipation.
@@ III.3 Landauer blowtorch effect emerges from localised electronic fluctuating and dissipating forces
Fig. 3(c) compares trajectory probability densities with and without the inclusion of electronic viscosity and noise; the figure also shows the spatial dependence of the effective temperature for $\chi=10$ meV at $V=1$ V. When only the electron force and environmental fluctuation–dissipation are included (orange dashed), the distribution is narrowly localised around the potential minimum. By contrast, the full treatment including electronic viscosity and noise (blue dashed) broadens the distribution, enabling access to higher-energy displacements even in the presence of a confining potential. This is a direct signature of the blowtorch effect: non-uniform dissipation and stochastic forces effectively create spatial variations in “temperature”, making escape pathways more accessible.